OpenStudy (gabylovesyou):

Which of the following is not true? A. 15 kilograms = 33.069 pounds B. 3 inches^3 = 49.161 centimeters^3 C. 30 kilometers = 18.642 miles D. 40 yards^2 = 63.576 meters^2

OpenStudy (anonymous):

1 kilogram = 2.20462262 pounds

OpenStudy (anonymous):

1 inches = 2.54 centimeters

OpenStudy (gabylovesyou):

A is true

OpenStudy (anonymous):

1 kilometer = 0.621371192 miles

OpenStudy (anonymous):

1 yds = 0.9144 meters

OpenStudy (anonymous):

have you studied dimensional analysis

OpenStudy (gabylovesyou):

mmm no ;s

OpenStudy (anonymous):

\[3(inch)(inch)(inch)*\frac{2.54cm}{1inch}*]\frac{2.54cm}{1inch}*\frac{2.54cm}{1inch}\]

OpenStudy (anonymous):

it's a technique used in science to get through different measurements... what you do is start out with the given and then use your measurement equalities as ratios

OpenStudy (anonymous):

you know from math thta when you have x/x , they cancel to get one. Dimensional analysis uses this same aspect but with the measurements like above

OpenStudy (anonymous):

now how would you check d using dimensional analysis

OpenStudy (april115):

i still dont get it

OpenStudy (anonymous):

what do you mean? this is not your question?!

OpenStudy (gabylovesyou):

lmaoooo

OpenStudy (anonymous):

anyways, do you understand gaby? start out with the given or one side of your equality

OpenStudy (gabylovesyou):

ok ill try :D

OpenStudy (anonymous):

you can use the equation editor if you'd like or when you get stuck i can help

OpenStudy (gabylovesyou):

ok thnks : D

OpenStudy (april115):

first of all i dont care if its not my ques. i was goin to help gabby. so when i saw the work that u did thats when i replied ''i dont get it''. understand?

OpenStudy (gabylovesyou):

can u do the second one so i see and ill try the one after?

OpenStudy (anonymous):

i'll start you off given:\[40yds^2\]

OpenStudy (gabylovesyou):

ok

OpenStudy (anonymous):

your unknown is : ? meters squared

OpenStudy (gabylovesyou):

right

OpenStudy (anonymous):

alright so how can you write the conversion as a ratio 1yd=.9144 meters

OpenStudy (gabylovesyou):

....................

OpenStudy (anonymous):

a ratio such as yds/meters or meters/yds

OpenStudy (mertsj):

\[40 yds^2\times\frac{.9144 m}{1 yd}\times\frac{.9144 m}{1yd}=33.44 m^2\]

OpenStudy (gabylovesyou):

so its D ! because its false!

OpenStudy (anonymous):

yes the ratios are the middle part of merts answer

OpenStudy (gabylovesyou):

ohh ok! Thank you soo much!~!

OpenStudy (anonymous):